OpenStudy (anonymous):

What is 7 1/5% of 465, rounded to the nearest tenth? a.) 32.5 b.) 35 c.) 33.5 d.) 34

OpenStudy (anonymous):

I know the answer is 33.5 but I need help setting up the equation.

OpenStudy (anonymous):

7 1/5%=0.072 0.072x 465= 33.5(nearest tenth)

OpenStudy (anonymous):

So it would be 34? @amstro

OpenStudy (anonymous):

nope 33.5

OpenStudy (anonymous):

For 7 1/5 what do I do exactly to get 0.072? (For the 1/5 part) @amstro

OpenStudy (anonymous):

1/5%=(1/5)/100

OpenStudy (anonymous):

@Zoobidoo

OpenStudy (anonymous):

ok i see. Thank you!